# Physics-Informed Time-Integrated DeepONet: Temporal Tangent Space Operator Learning for High-Accuracy Inference

### v2

@@ 3.2 One-dimensional Burgers’ Equation
The dataset is generated by solving Eq. 10 on a periodic spatial domain $x\in[0,1]$. The spatial domain is discretized using $s=4,096$ grid points, while outputs are captured at $n_{\text{steps}}+1=101$ evenly spaced time steps over $t\in[0,1]$. The ICs $u_{0}(x)$ are sampled from stationary Gaussian random fields (GRFs) with covariance kernel $C=\sigma^{2}(-\Delta+\tau^{2}I)^{-\gamma}$. For GRF generation, the Fourier coefficients are drawn from $\mathcal{N}(0,\sigma^{2})$ and scaled appropriately to enforce the specified covariance structure, yielding smooth, periodic initial profiles. The Burgers’ equation is advanced in time using an operator-splitting spectral method implemented with the spin function from Chebfun [chebfun], which combines linear evolution via viscosity $\nu u_{xx}$ with nonlinear convection $uu_{x}$. The solution $u(t,x)$ is represented as a Chebyshev interpolant at each time step and stored. The corresponding time derivative $\partial_{t}u(t,x)$ is independently computed at each output time by substituting the solution into the original Burgers’ equation. Consistency checks are performed to ensure the accuracy of $\partial_{t}u(t,x)$ using both finite differences and the governing PDE. The dataset consists of $N=2,500$ realizations of $u_{0}(x)$, where each realization evolves under the Burgers’ dynamics and is evaluated on a reduced grid of $n=101$ spatial points.
@@
Figure 9: 1D Burgers’ equation: Comparison of different time integration (PITI) methods using explicit Euler, fourth-order Runge-Kutta (RK4) and second-order Adam-Bashforth-Moultons (ABM2) with mean relative $\mathcal{L}_{2}$ error on the left and predicted residuals on the right over all test examples.
@@
The full trajectories as well as the residuals obtained during inference are shown in Fig. 13, with full field plots available in the Fig. 10. The residuals for different time steps (Fig. 13) as well as different methods (Fig. 9) indicate a regime around the middle of the testing domain that was learned less accurately compared to states before and after. This is discernible due to an increase in the mean residual over all test examples, followed by a reduction back to smaller residual values. One could exploit this to adaptively resample profiles from this region of increased residuals in order to reduce the overall error. The MAE between the squared difference and residual for explicit Euler with $\Delta t=0.01$ is $2.63\mathrm{e}\text{-}{-5}$ with Pearson’s correlation coefficient $\rho=0.9996$, once again suggesting that this correlation can be leveraged to obtain higher quality predictions. The last study for this example includes the already displayed classic form in a hybrid setting against the special form for time-evolution PDEs (mean relative $\mathcal{L}_{2}$ error $3.62\mathrm{e}\text{-}2$) alongside the classic form without additional data (mean relative $\mathcal{L}_{2}$ error $2.07\mathrm{e}\text{-}1$). Thus, the chosen setup outperforms the data-free setting by more than an order of magnitude while still halving the error compared to the special case, thus indicating better performance on more complex equations.
@@ 3.3 Two-dimensional Allen-Cahn Equation
Furthermore, the chosen time integration scheme during inference has negligible influence on the errors (Tab. 1 and Fig. 16). An example including error comparison for the methods is shown in Fig. 14(b), demonstrating the struggle of AR to capture the dynamics even within the training domain, and FR starting to slightly diverge in extrapolation with relatively low dynamics in the remaining field. In contrast, PITI exhibits comparably more gradual divergence, yet its residual continues to grow in the later states, thus hinting at the potential to improve sampling of states to sufficiently represent the underlying dynamics (Fig. 16). Once again, residual and squared difference exhibit near perfect correlation, with a MAE of $1.84\mathrm{e}\text{-}3$ and Pearson’s correlation coefficient $\rho=0.997$ for explicit Euler with $\Delta t=0.01$.
@@
Figure 16: 2D Allen-Cahn equation: Comparison of different time integration (PITI) methods using explicit Euler, fourth-order Runge-Kutta (RK4) and second-order Adam-Bashforth-Moultons (ABM2) with mean relative $\mathcal{L}_{2}$ error on the left and predicted residuals on the right over all test examples.
@@ 3.4 Ablation Studies
Table 4: Seed sensitivity analysis for the final model configurations. Each method (PITI, AR, FR) was trained with five different random seeds. For each run, the relative $\mathcal{L}_{2}$ error was computed over all test samples and summarized by its mean, minimum, and maximum at two time points corresponding to the end of the training domain and the end of the testing domain for each problem (Heat 1D: $t=0.5$ and $t=5.0$; Burgers 1D: $t=0.5$ and $t=1.0$; Allen-Cahn 2D: $t=0.5$ and $t=1.0$). Table entries report the overall mean and between-seed standard deviation across the five runs, formatted as $\mu\pm\sigma$.
Heat 1D Burgers 1D Allen-Cahn 2D
$t=0.5$ $t=5.0$ $t=0.5$ $t=1.0$ $t=0.5$ $t=1.0$
Mean
PITI $3.5\mathrm{e}\text{-}2\pm 3.9\mathrm{e}\text{-}3$ $2.2\mathrm{e}\text{-}1\pm 3.4\mathrm{e}\text{-}2$ $1.1\mathrm{e}\text{-}2\pm 6.2\mathrm{e}\text{-}4$ $1.8\mathrm{e}\text{-}2\pm 1.8\mathrm{e}\text{-}3$ $7.9\mathrm{e}\text{-}2\pm 1.4\mathrm{e}\text{-}3$ $1.2\mathrm{e}\text{-}1\pm 2.6\mathrm{e}\text{-}3$
AR $2.2\mathrm{e}\text{-}1\pm 3.3\mathrm{e}\text{-}2$ $1.5\mathrm{e}\text{-}0\pm 5.4\mathrm{e}\text{-}1$ $2.4\mathrm{e}\text{-}1\pm 8.7\mathrm{e}\text{-}2$ $6.0\mathrm{e}\text{-}1\pm 2.1\mathrm{e}\text{-}1$ $9.9\mathrm{e}\text{-}1\pm 7.2\mathrm{e}\text{-}2$ $1.1\mathrm{e}0\pm 7.0\mathrm{e}\text{-}2$
FR $5.7\mathrm{e}\text{-}3\pm 3.2\mathrm{e}\text{-}4$ $1.3\mathrm{e}0\pm 1.4\mathrm{e}\text{-}1$ $3.5\mathrm{e}\text{-}2\pm 2.0\mathrm{e}\text{-}3$ $1.2\mathrm{e}\text{-}1\pm 2.1\mathrm{e}\text{-}2$ $1.3\mathrm{e}\text{-}1\pm 2.3\mathrm{e}\text{-}2$ $2.1\mathrm{e}\text{-}1\pm 1.8\mathrm{e}\text{-}2$
Min
PITI $3.8\mathrm{e}\text{-}3\pm 4.2\mathrm{e}\text{-}4$ $2.8\mathrm{e}\text{-}2\pm 1.7\mathrm{e}\text{-}3$ $2.5\mathrm{e}\text{-}3\pm 3.9\mathrm{e}\text{-}4$ $4.4\mathrm{e}\text{-}3\pm 5.9\mathrm{e}\text{-}4$ $5.3\mathrm{e}\text{-}2\pm 1.5\mathrm{e}\text{-}3$ $7.7\mathrm{e}\text{-}2\pm 3.6\mathrm{e}\text{-}3$
AR $3.2\mathrm{e}\text{-}2\pm 8.3\mathrm{e}\text{-}3$ $1.3\mathrm{e}\text{-}1\pm 5.9\mathrm{e}\text{-}2$ $4.1\mathrm{e}\text{-}2\pm 7.0\mathrm{e}\text{-}3\mathdollar$ $1.0\mathrm{e}\text{-}1\pm 3.3\mathrm{e}\text{-}2$ $1.8\mathrm{e}\text{-}1\pm 3.6\mathrm{e}\text{-}2$ $1.9\mathrm{e}\text{-}1\pm 4.3\mathrm{e}\text{-}2$
FR $1.4\mathrm{e}\text{-}3\pm 3.4\mathrm{e}\text{-}4$ $6.3\mathrm{e}\text{-}1\pm 1.3\mathrm{e}\text{-}1$ $7.0\mathrm{e}\text{-}3\pm 7.2\mathrm{e}\text{-}4$ $3.9\mathrm{e}\text{-}2\pm 1.0\mathrm{e}\text{-}2$ $4.0\mathrm{e}\text{-}2\pm 7.1\mathrm{e}\text{-}3$ $7.1\mathrm{e}\text{-}2\pm 1.3\mathrm{e}\text{-}2$
Max
PITI $1.7\mathrm{e}\text{-}1\pm 3.4\mathrm{e}\text{-}2$ $2.5\mathrm{e}\text{-}0\pm 6.6\mathrm{e}\text{-}1$ $1.9\mathrm{e}\text{-}1\pm 3.7\mathrm{e}\text{-}2$ $2.5\mathrm{e}\text{-}1\pm 7.1\mathrm{e}\text{-}2$ $1.1\mathrm{e}\text{-}1\pm 6.0\mathrm{e}\text{-}3$ $1.8\mathrm{e}\text{-}1\pm 1.2\mathrm{e}\text{-}2$
AR $9.7\mathrm{e}\text{-}1\pm 2.1\mathrm{e}\text{-}1$ $3.2\mathrm{e}1\pm 1.7\mathrm{e}1$ $1.8\mathrm{e}0\pm 1.0\mathrm{e}0$ $5.2\mathrm{e}0\pm 3.5\mathrm{e}0$ $3.3\mathrm{e}0\pm 4.1\mathrm{e}\text{-}1$ $2.9\mathrm{e}0\pm 3.8\mathrm{e}\text{-}1$
FR $2.9\mathrm{e}\text{-}2\pm 5.9\mathrm{e}\text{-}3$ $2.1\mathrm{e}1\pm 2.5\mathrm{e}0$ $2.5\mathrm{e}\text{-}1\pm 3.7\mathrm{e}\text{-}2$ $5.3\mathrm{e}\text{-}1\pm 9.3\mathrm{e}\text{-}2$ $3.4\mathrm{e}\text{-}1\pm 6.8\mathrm{e}\text{-}2$ $4.7\mathrm{e}\text{-}1\pm 6.6\mathrm{e}\text{-}2$

### v3

@@ 3.2 One-dimensional Burgers’ Equation
The dataset is generated by solving Eq. 13 on a periodic spatial domain $x\in[0,1]$. The spatial domain is discretized using $s=4,096$ grid points, while outputs are captured at $n_{\text{steps}}+1=101$ evenly spaced time steps over $t\in[0,1]$. The ICs $u_{0}(x)$ are sampled from stationary Gaussian random fields (GRFs) with covariance kernel $C=\sigma^{2}(-\Delta+\tau^{2}I)^{-\gamma}$. For GRF generation, the Fourier coefficients are drawn from $\mathcal{N}(0,\sigma^{2})$ and scaled appropriately to enforce the specified covariance structure, yielding smooth, periodic initial profiles. The Burgers’ equation is advanced in time using an operator-splitting spectral method implemented with the spin function from Chebfun Driscoll et al. [2014], which combines linear evolution via viscosity $\nu u_{xx}$ with nonlinear convection $uu_{x}$. The solution $u(t,x)$ is represented as a Chebyshev interpolant at each time step and stored. The corresponding time derivative $\partial_{t}u(t,x)$ is independently computed at each output time by substituting the solution into the original Burgers’ equation. Consistency checks are performed to ensure the accuracy of $\partial_{t}u(t,x)$ using both finite differences and the governing PDE. The dataset consists of $N=2,500$ realizations of $u_{0}(x)$, where each realization evolves under the Burgers’ dynamics and is evaluated on a reduced grid of $n=101$ spatial points.
@@
Figure 9: 1D Burgers’ equation: Comparison of different time integration (PITI) methods using explicit Euler, fourth-order Runge-Kutta (RK4), and second-order Adam-Bashforth-Moultons (ABM2) with mean relative $\mathcal{L}_{2}$ error on the left and predicted residuals on the right over all test examples. Note that in both subplots several curves overlap almost exactly and therefore appear as a single trace.
@@
The full trajectories as well as the residuals obtained during inference are shown in Fig. 14, with full field plots available in the Fig. 10. The residuals for different time steps (Fig. 14) as well as different methods (Fig. 9) indicate a regime around the middle of the testing domain that was learned less accurately compared to states before and after. This is discernible due to an increase in the mean residual over all test examples, followed by a reduction back to smaller residual values. One could exploit this to adaptively resample profiles from this region of increased residuals in order to reduce the overall error. The MAE between the squared difference and residual for explicit Euler with $\Delta t=0.01$ is $2.63\mathrm{e}\text{-}{-5}$ with Pearson’s correlation coefficient $\rho=0.9996$, once again suggesting that this correlation can be leveraged to obtain higher quality predictions. The last study for this example includes the already displayed classic form in a hybrid setting against the special form for explicit time-evolution PDEs (mean relative $\mathcal{L}_{2}$ error $3.62\mathrm{e}\text{-}2$) alongside the classic form without additional data (mean relative $\mathcal{L}_{2}$ error $2.07\mathrm{e}\text{-}1$). As shown in Fig. 12, the special formulation (cf. Section 2.2.1) does not provide an accuracy benefit for the more nonlinear or advection-dominated Burgers’ equation, and the vanilla formulation remains more robust. Thus, the chosen setup outperforms the data-free setting by more than an order of magnitude while still halving the error compared to the special case, thus indicating better performance on more complex equations.
Since Burgers’ equation has been extensively studied in the literature, we provide two additional comparisons to established baselines. First, we consider physics-constrained deep auto-regressive networks (PCAR, Geneva and Zabaras [2020]). We adopt the hyperparameter recommendations of Geneva and Zabaras [2020], but modify the architecture to match our setting by using a single input slice. We benchmark this adapted PCAR model against a physics-informed DeepONet AR approach and PITI-DeepONet with an explicit Euler time integrator. All three methods are trained under identical conditions with purely physics-informed training using the same data and the same hyperparameters specified above. Test-set results are summarized in Fig. 13. Overall, PCAR performs worse than DeepONet-based AR and PITI. While PCAR achieves a final relative $\mathcal{L}_{2}$ error of $2.95\mathrm{e}\text{-}2$ on the held-out validation set (computed over all slices across the full time horizon $t\in[0,1]$), its rollout yields a mean relative $\mathcal{L}_{2}$ error of $2.41$ on the final slice of the test domain (min: $1.03$, max: $3.66$). In contrast, AR attains $1.43$ (min: $2.78\mathrm{e}\text{-}1$, max: $1.69$), and PITI achieves $2.06\mathrm{e}\text{-}1$ (min: $2.52\mathrm{e}\text{-}2$, max: $2.16$).
Figure 13: 1D Burgers’ equation: Time evolution of the relative $\mathcal{L}_{2}$ error for PCAR, AR, and PITI with explicit Euler in a purely physics-informed setting. The PITI ansatz reduces error growth relative to both auto-regressive rollouts.
Second, we compare against the reported MSE in PINODE Sholokhov et al. [2023].PINODE samples initial conditions from a finite harmonic family and considers 128 spatial points for $x\in[-\pi,\pi]$ and $t\in[0,2]$ with $\Delta t=0.1$ in training. In contrast, our setup samples initial conditions from a Gaussian random field, uses $x\in[0,1]$ with 101 spatial points, and considers $t\in[0,1]$ with 101 time steps. While this differs from our setup, PINODE reports MSE values for unrolling beyond the training domain; hence, we include this as a reference. We reference the reported mean MSE in the first post-training period as $7\mathrm{e}\text{-}2$ (physics-only) and $5\mathrm{e}\text{-}3$ (hybrid) as as no official PINODE code is available for a direct re-implementation. While recognizing that the underlying data production and domains differ between the two studies, PITI with explicit Euler achieves $1.1\mathrm{e}\text{-}3$ (physics-informed) and $3.3\mathrm{e}\text{-}5$ (hybrid) in the post-training phase under the same period-based MSE, suggesting significantly lower error in this situation.
@@ 3.3 Two-dimensional Allen-Cahn Equation
Furthermore, the chosen time integration scheme during inference has negligible influence on the errors (Tab. 1 and Fig. 17). An example including error comparison for the methods is shown in Fig. 15(b), demonstrating the struggle of AR to capture the dynamics even within the training domain, and FR starting to slightly diverge in extrapolation with relatively low dynamics in the remaining field. In contrast, PITI exhibits comparably more gradual divergence, yet its residual continues to grow in the later states, thus hinting at the potential to improve sampling of states to sufficiently represent the underlying dynamics (Fig. 17). Once again, residual and squared difference exhibit near perfect correlation, with an MAE of $1.84\mathrm{e}\text{-}3$ and Pearson’s correlation coefficient $\rho=0.997$ for explicit Euler with $\Delta t=0.01$.
@@
Figure 17: 2D Allen-Cahn equation: Comparison of different time integration (PITI) methods using explicit Euler, fourth-order Runge-Kutta (RK4) and second-order Adam-Bashforth-Moultons (ABM2) with mean relative $\mathcal{L}_{2}$ error on the left and predicted residuals on the right over all test examples. Note that in both subplots several curves overlap almost exactly and therefore appear as a single trace.
3.4 One-dimensional Kuramoto-Sivashinsky Equation
Finally, to demonstrate the applicability of PITI to more chaotic systems, we consider the Kuramoto-Sivashinsky (KS) equation. The equation reads as:
$\frac{\partial u}{\partial t}=-\frac{\partial^{2}u}{\partial x^{2}}-\frac{\partial^{4}u}{\partial x^{4}}-u\frac{\partial u}{\partial x},\quad(t,x)\in[0,\infty)\times[0,22]$ (15)
with periodic BCs $u(t,0)=u(t,22)$ and $\tfrac{\partial u}{\partial x}(t,0)=\tfrac{\partial u}{\partial x}(t,22)$. The initial condition $u(0,x)=f(x)$ was generated by drawing a random periodic profile $u_{0}(x)$ from i.i.d. Gaussian coefficients and evolve the PDE using the Fourier spectral spin solver in Chebfun, with an internal step-size cap $\Delta t_{\max}=10^{-3}$. After a burn-in of $t_{\mathrm{burn}}=23$ time units, which mitigates startup transients, we set $f(x)=u(t_{\mathrm{burn}},x)$ as effective IC thus yielding $f$ representative of the KS long-time statistically stationary regime. From this state, we record $101$ uniformly spaced snapshots over $T=10$ time units and evaluate the solution on a reduced grid of $n=128$ spatial points. Time derivatives $\partial_{t}u(t,x)$ are stored alongside $u(t,x)$ (computed from the KS right-hand side) and cross-checked on the test set using finite differences. In total, the dataset consists of $N_{\mathrm{train}}=2000$ training and $N_{\mathrm{test}}=500$ test trajectories.
We implement the hybrid versions for FR, AR, and PITI with 1,200 IC profiles sampled at $t=0$ for purely physics-informed training (TI at $t=0$, AR with $\rm{dt}=0.01$, FR on $t\in[0,5.0]$) and additionally 1,200 profiles sampled from 400 trajectories at $t=\{0,2.5,5.0\}$ used for both physics-informed and data-driven training. This is the same setup as for Burgers’ equation, but adapted to the longer time period for KS. Apart from that, we restrict our baselines to explicit Euler and additionally include backward Euler for stability, while omitting ABM2 and RK4 as inference schemes for PITI. Fig. 2(d) depicts a comparison of the relative $\mathcal{L}_{2}$ error over time, showcasing the inability of FR to capture any dynamics with the given information, while AR deteriorates about a magnitude faster than PITI, which is also discernible in Fig 18.
Figure 18: 1D Kuramoto-Sivashinsky equation: Reference and difference of predictions to reference using time integration with implicit Euler.
While the resulting relative $\mathcal{L}_{2}$ errors were comparable, the reconstruction errors during inference were lower for implicit Euler as given in Fig. 19. This is particularly interesting as the underlying operator is identical, and the numerical update seems to keep the state within the learned domain for the A-stable implicit Euler. Once again, linear correlation between reconstruction error and error in the time derivative is almost perfect with $\rho=0.974$. This slightly worse value compared to the other three methods seems to stem from the more pronounced error on the field value with a relative $\mathcal{L}_{2}$ error of $2.5\mathrm{e}\text{-}1$ on the held-out validation data compared to the error for the time derivative with $6.8\mathrm{e}\text{-}2$ at the end of training.
Figure 19: 1D Kuramoto-Sivashinsky equation: Reference and difference of predictions to reference using time integration with implicit Euler. Only the first and last residual in each iteration for IE is shown.
These results for KS showcase the strength of PITI to work for chaotic equations with limited data, where FR fails to capture any underlying dynamics even on the training data within the training domain with a relative $\mathcal{L_{2}}$ errors of 0.69. AR was able to capture this dynamic with relative $\mathcal{L}_{2}$ errors of $1.63\mathrm{e}\text{-}2$ on the train and $1.32\mathrm{e}\text{-}2$ on the validation data at the end of training, but during inference, the predicted trajectory rapidly diverges from the reference solution. Even though the employed time-integration schemes are not standard choices for the stiff, chaotic Kuramoto–Sivashinsky equation, they still deliver markedly better performance than the reference methods using operator learning in a hybrid physics-informed and limited data-driven setup.
@@ 3.5 Ablation Studies
To further characterize the robustness of the results, we investigate how random initialization, training-set size, and the inference time step size influence the reported performance.
@@
Table 4: Seed sensitivity analysis for the final model configurations. Each method (PITI, AR, FR) was trained with five different random seeds. For each run, the relative $\mathcal{L}_{2}$ error was computed over all test samples and summarized by its mean, minimum, and maximum at two time points corresponding to the end of the training domain and the end of the testing domain for each problem (Heat 1D: $t=0.5$ and $t=5.0$; Burgers 1D: $t=0.5$ and $t=1.0$; Allen-Cahn 2D: $t=0.5$ and $t=1.0$; Kuramoto-Sivashinsky (KS) 1D: $t=5.0$ and $t=10.0$). Table entries report the overall mean and between-seed standard deviation across the five runs, formatted as $\mu\pm\sigma$.
Heat 1D Burgers 1D Allen-Cahn 2D KS 1D
$t=0.5$ $t=5.0$ $t=0.5$ $t=1.0$ $t=0.5$ $t=1.0$ $t=5.0$ $t=10.0$
Mean PITI $3.5\mathrm{e}\text{-}2\pm 3.9\mathrm{e}\text{-}3$ $2.2\mathrm{e}\text{-}1\pm 3.4\mathrm{e}\text{-}2$ $1.1\mathrm{e}\text{-}2\pm 6.2\mathrm{e}\text{-}4$ $1.8\mathrm{e}\text{-}2\pm 1.8\mathrm{e}\text{-}3$ $7.9\mathrm{e}\text{-}2\pm 1.4\mathrm{e}\text{-}3$ $2\mathrm{e}\text{-}1\pm 2.6\mathrm{e}\text{-}3 1.1$$1\mathrm{e}\text{-}1\pm 9.5\mathrm{e}\text{-}3$ $4.2\mathrm{e}\text{-}1\pm 3.4\mathrm{e}\text{-}2$
AR $2.2\mathrm{e}\text{-}1\pm 3.3\mathrm{e}\text{-}2$ $1.5\mathrm{e}0\pm 5.4\mathrm{e}\text{-}1$ $2.4\mathrm{e}\text{-}1\pm 8.7\mathrm{e}\text{-}2$ $6.0\mathrm{e}\text{-}1\pm 2.1\mathrm{e}\text{-}1$ $9.9\mathrm{e}\text{-}1\pm 7.2\mathrm{e}\text{-}2$ $1.1\mathrm{e}0\pm 7.0\mathrm{e}\text{-}2$ $8.8\mathrm{e}\text{-}1\pm 2.8\mathrm{e}\text{-}1$ $1.2\mathrm{e}0\pm 4.5\mathrm{e}\text{-}2$
FR $5.7\mathrm{e}\text{-}3\pm 3.2\mathrm{e}\text{-}4$ $1.3\mathrm{e}0\pm 1.4\mathrm{e}\text{-}1$ $3.5\mathrm{e}\text{-}2\pm 2.0\mathrm{e}\text{-}3$ $1.2\mathrm{e}\text{-}1\pm 2.1\mathrm{e}\text{-}2$ $1.3\mathrm{e}\text{-}1\pm 2.3\mathrm{e}\text{-}2$ $2.1\mathrm{e}\text{-}1\pm 1.8\mathrm{e}\text{-}2$ $6.9\mathrm{e}\text{-}1\pm 6.1\mathrm{e}\text{-}2$ $1.0\mathrm{e}0\pm 7.5\mathrm{e}\text{-}2$
Min PITI $3.8\mathrm{e}\text{-}3\pm 4.2\mathrm{e}\text{-}4$ $2.8\mathrm{e}\text{-}2\pm 1.7\mathrm{e}\text{-}3$ $2.5\mathrm{e}\text{-}3\pm 3.9\mathrm{e}\text{-}4$ $4.4\mathrm{e}\text{-}3\pm 5.9\mathrm{e}\text{-}4$ $5.3\mathrm{e}\text{-}2\pm 1.5\mathrm{e}\text{-}3$ $7.7\mathrm{e}\text{-}2\pm 3.6\mathrm{e}\text{-}3$ $2.9\mathrm{e}\text{-}2\pm 3.8\mathrm{e}\text{-}3$ $7.5\mathrm{e}\text{-}2\pm 4.7\mathrm{e}\text{-}3$
AR $3.2\mathrm{e}\text{-}2\pm 8.3\mathrm{e}\text{-}3$ $1.3\mathrm{e}\text{-}1\pm 5.9\mathrm{e}\text{-}2$ $4.1\mathrm{e}\text{-}2\pm 7.0\mathrm{e}\text{-}3$ $1.0\mathrm{e}\text{-}1\pm 3.3\mathrm{e}\text{-}2$ $1.8\mathrm{e}\text{-}1\pm 3.6\mathrm{e}\text{-}2$ $1.9\mathrm{e}\text{-}1\pm 4.3\mathrm{e}\text{-}2$ $2.2\mathrm{e}\text{-}1\pm 1.2\mathrm{e}\text{-}1$ $3.1\mathrm{e}\text{-}1\pm 6.7\mathrm{e}\text{-}2$
FR $1.4\mathrm{e}\text{-}3\pm 3.4\mathrm{e}\text{-}4$ $6.3\mathrm{e}\text{-}1\pm 1.3\mathrm{e}\text{-}1$ $7.0\mathrm{e}\text{-}3\pm 7.2\mathrm{e}\text{-}4$ $3.9\mathrm{e}\text{-}2\pm 1.0\mathrm{e}\text{-}2$ $4.0\mathrm{e}\text{-}2\pm 7.1\mathrm{e}\text{-}3$ $7.1\mathrm{e}\text{-}2\pm 1.3\mathrm{e}\text{-}2$ $1.2\mathrm{e}\text{-}1\pm 5.2\mathrm{e}\text{-}2$ $4.9\mathrm{e}\text{-}1\pm 1.5\mathrm{e}\text{-}1$
Max PITI $1.7\mathrm{e}\text{-}1\pm 3.4\mathrm{e}\text{-}2$ $2.5\mathrm{e}0\pm 6.6\mathrm{e}\text{-}1$ $1.9\mathrm{e}\text{-}1\pm 3.7\mathrm{e}\text{-}2$ $2.5\mathrm{e}\text{-}1\pm 7.1\mathrm{e}\text{-}2$ $1.1\mathrm{e}\text{-}1\pm 6.0\mathrm{e}\text{-}3$ $1.8\mathrm{e}\text{-}1\pm 1.2\mathrm{e}\text{-}2$ $4.0\mathrm{e}\text{-}1\pm 6.3\mathrm{e}\text{-}2$ $1.3\mathrm{e}0\pm 1.1\mathrm{e}\text{-}1$
AR $9.7\mathrm{e}\text{-}1\pm 2.1\mathrm{e}\text{-}1$ $3.2\mathrm{e}1\pm 1.7\mathrm{e}1$ $1.8\mathrm{e}0\pm 1.0\mathrm{e}0$ $5.2\mathrm{e}0\pm 3.5\mathrm{e}0$ $3.3\mathrm{e}0\pm 4.1\mathrm{e}\text{-}1$ $2.9\mathrm{e}0\pm 3.8\mathrm{e}\text{-}1$ $2.1\mathrm{e}0\pm 2.5\mathrm{e}\text{-}1$ $2.4\mathrm{e}0\pm 3.0\mathrm{e}\text{-}1$
FR $2.9\mathrm{e}\text{-}2\pm 5.9\mathrm{e}\text{-}3$ $2.1\mathrm{e}1\pm 2.5\mathrm{e}0$ $2.5\mathrm{e}\text{-}1\pm 3.7\mathrm{e}\text{-}2$ $5.3\mathrm{e}\text{-}1\pm 9.3\mathrm{e}\text{-}2$ $3.4\mathrm{e}\text{-}1\pm 6.8\mathrm{e}\text{-}2$ $4.7\mathrm{e}\text{-}1\pm 6.6\mathrm{e}\text{-}2$ $1.7\mathrm{e}0\pm 6.1\mathrm{e}\text{-}1$ $2.2\mathrm{e}0\pm 2.6\mathrm{e}\text{-}1$
3.5.1 Random Seed